OpenStudy (anonymous):

what is the answer to the following equation? y-3x=-4 and y=2x-1

OpenStudy (mathteacher1729):

You are looking for "The intersection of two lines". If you would like to do so algebraically: \[2x-1=3x-4\] Now subtract 2x from both sides \[-1=x-4\] Now add 4 to both sides \[3=x\] Now substitute x=3 into y = 2x-1 \[y=2(3)-1\] So y=5. You should see this by graphing as well. On your calculator: \[y_1=3x-4\] \[y_2=2x-1\] And see that they intersect at the point (3,5). Hope this helps.

OpenStudy (anonymous):

what is the answer to the following equation? 3x+6y=0 and y=4x+9

OpenStudy (mathteacher1729):

Mouse -- did you understand my answer to your original question?

OpenStudy (anonymous):

not really

OpenStudy (mathteacher1729):

Mouse -- were there any parts of the answer that made sense?

OpenStudy (anonymous):

yes like the first part of it but then i got confused.

OpenStudy (mathteacher1729):

Ok. That's a good start! :) If you can let me know exactly where it started to get confusing... I can help you.

OpenStudy (anonymous):

its starts getting confusing when you have to substitute.

OpenStudy (mathteacher1729):

So you understand how I got x=3? Once we have x = 3, then we can take that value of x (which is half the answer, we're looking for an x and a y) and insert it into either of the original equations. That lets us solve for y.

OpenStudy (anonymous):

oh. ok. i get it now. thanks

OpenStudy (mathteacher1729):

Awesome!